Hannah o ch s a soccer lea in Bowling Green. T or ow g hey have a r Supp se the e s a 44% chance ha he earn wil use. Regardless of it her er e m oses, H rinah uns a achine a aren wil complain about something o her after he game. From prev s seasons, Hannah nows hat he probabi y the her earn oses ㎸n parent complains is 319 Organ ze and be the g verrform on rs be ore star ng W it ou he probability stame for each question before calculating anything.) Round al probailities to four decimas) a Suppose Hannอhs soccer team loses. What is the probability that parent complains? ti If a parent complains after the game, what is the probahility that Hannahs team last? c What is the probability that Hannahs team wll win tomorrow night? d. Suppose Hannahs soocer tcam loscs. What is the probability that a parent docs not complain?
Answer #1

L - team will lose

C -complain

P(L) = 0.44

P(C) = 0.51

P(C and L) = 0.31

a)

P(C|L) = P( C and L)/P(L)

= 0.31/0.44

= 0.704545

b)

P(L|C) = P(C an L)/P(C) = 0.31/0.51

= 0.6078431

c)

P(Win) = 1- P(L) = 1 - 0.44 = 0.56

d)

P(C'| L) = P(C' and L) /P(L)

= ( P(L) - P(L and C)) / 0.44

= ( 0.44- 0.31)/ 0.44

= 0.295454

Please rate